If the equation of the circle 4x^2+24x+4y^2-32y=44 can be written as (x-a)^2+(y-b)^2=r^2, then a+b+r=

A) 12
B) 13
C) -1
D) 7
E) 5

Answer by rothauserc(4718)

4x^2 +24x +4y^2 -32y = 44
divide both sides of = by 4
x^2 +6x +y^2 -8y = 11
complete the square
x^2 +6x +9 +y^2 -8y + 16 = 11 +9 +16
(x+3)^2 +(y-4)^2 = 36, therefore
a = -3, b = 4, r = 6, then
-3 +4 +6 = 7
answer is D